The pair of values below is from a direct variation. Find the missing number.
(9,5) and (x,3)


Answer by stanbon(75887)

The pair of values below is from a direct variation. Find the missing number.
(9,5) and (x,3)
-----
x/3 = 9/5
x = 27/5
x = 5.4
--------------
OR
x/9 = 3/5
x = 27/5
x = 5.4
---------------
Cheers,
Stan H.
----------------